# Limited-Range Multilinear Off-Diagonal Extrapolation and Weighted Transference Principle

### v1

@@ 1 Introduction
As a matter of fact, this symmetry also shows its benefits for the extrapolation results. The theory of $L^{p}$-extrapolation has a long history starting with the works by García-Cuerva and Rubio de Francia [7, 16], who realized that boundedness of an operator $T$ on all Muckenhoupt weighted $L^{p}$-spaces implies boundedness on all Muckenhoupt weighted $L^{q}$-spaces for the range $q\in(1,\infty)$. They also showed that vector-valued estimates can be extracted from scalar-valued weighted estimates. Later, Buckley [2] found quantitative bounds for the maximal operator in Muckenhoupt weighted spaces in terms of the weight characteristic. Restricting the weight classes for the initial spaces of extrapolation, Auscher and Martell proved first versions of limited-range extrapolation in [1]. For a detailed history of the theory, we refer to Nieraeth [12], where a comprehensive theory for multilinear extrapolation is established. In particular, in the multilinear case $T:\prod_{j=1}^{m}L^{p_{j}}_{{w}_{j}}\to L^{p}_{{w}}$ with $\frac{1}{p}=\sum_{j=1}^{m}\frac{1}{p_{j}}$ some of the $p_{j}$’s (but not all of them) are allowed to be infinity in the target spaces of the extrapolation. In [10], the authors provided conditions that ensure extrapolation towards $p=\infty$. In the linear case, they also showed that this is possible in the off-diagonal setting, that is when $p\neq p_{1}$. Recently, Nieraeth [13] found extrapolation results from Lebesgue spaces over $\sigma$-finite measure to general quasi-Banach function spaces, which includes mixed-norm estimates. Our contribution to the theory of extrapolation in weighted spaces is as follows:
@@
    We extend the results of [10]: We include the range $(0,1]$ for all initial and target spaces of the extrapolation. We show off-diagonal estimates even in the multilinear case and allow for mixed $L^{\mathbf{p}}$-norms and two-weight estimates. We provide explicit bounds, which are sharp in all scalar-valued cases considered in [10] (where no explicit bounds are provided). In comparison to [10] our proofs are significantly shorter and leaner. This is achieved by systematically exploiting the symmetry of the weight classes. In fact, Theorem 2.6 contains Theorem 2.3 in [10], where the authors are forced to consider three different cases, by specialising to the case $p_{0}\in[1,\infty]$, $p\in(1,\infty]$ and $r_{0}:=p_{0}^{\prime}$, $r=p^{\prime}$. We warn the reader that our $s$ is called $r$ in [10]. Finally, we work in the setting of general $\sigma$-finite measure spaces as in [13] instead of $\mathbb{R}^{d}$.
@@
We fix $(\Omega,\mu)$ together with a basis of sets $U$ for the rest of this section and write $L^{0}(\Omega;X)$ for the space of strongly measurable functions, where $X$ is a quasi-Banach space. In the scalar-valued case, we simply write $L^{0}(\Omega)$.
@@
Note that ${\mathcal{M}}f$ is the countable supremum of measurable functions and hence measurable. Moreover, for $f\in L^{1}(\Omega)$ we have ${\mathcal{M}}f>0$ whenever $f\neq 0$, see [13, Proposition 2.16].
@@
    $\|{\mathcal{M}}\|_{L^{t}_{{w}}(\Omega)\to L^{t}_{{w}}(\Omega)}\leq c(t)[{w}]_{t}^{t^{\prime}}\quad\text{for all $t\in(1,\infty]$ and weights ${w}$}.$
@@
    As observed in [13, Remark 4.8], the formulation in Theorem 1.2 implies a result in terms of extrapolation pairs. Indeed, if
@@
    then the choice $V:=\mathcal{F}$, $S(f,g):=g$, and $T(f,g):=f$ shows that a bound $\|g\|_{L^{q_{0}}_{{w}_{0}}(\Omega)}\leq\phi([{w}]_{(s_{0},r_{0})})\|f\|_{L^{p_{0}}_{{w}_{0}}(\Omega)}$ for all $(f,g)\in\mathcal{F}$ and weights ${w}_{0}$ with ${w}_{0}]_{(s_{0},r_{0})}<\infty$ extrapolates to a bound $\|g\|_{L^{q}_{{w}}(\Omega)}\leq\phi_{s_{0},r_{0},s}([{w}]_{(s,r)})\|f\|_{L^{p}_{{w}}(\Omega)}$ for all $(f,g)\in\mathcal{F}$ and weights w{w} with ${w}]_{(s,r)}<\infty$.
@@
    We can deduce an (unweighted) bound $L^{1}(\mathbb{R}^{d})\to L^{\infty}(\mathbb{R}^{d})$ whenever we already know a bound $L^{\frac{2}{3}}_{{w}_{0}}(\mathbb{R}^{d})\to L^{2}_{{w}_{0}}(\mathbb{R}^{d})$ for all ${w}]_{1}<\infty$. Indeed, setting
@@
    we obtain this assertion as a particular instance of Theorem 1.1 since $1]_{2}=1<\infty$ and since $\|{\mathcal{M}}\|_{L^{t}_{{w}}(\mathbb{R}^{d})\to L^{t}_{{w}}(\mathbb{R}^{d})}\leq c(t)[{w}]_{t}^{t^{\prime}}$ for all $t\in(1,\infty]$ and weights w{w}. This shows that by detaching the exponents for the weight classes from the exponents for the initial and target spaces for the extrapolation, Theorem 1.1 can cover situations that can not be treated by [10, Theorem 2.3], where $p$ has to be strictly larger than $1$.
@@ 2 Limited-Range Multilinear Off-Diagonal Extrapolation
$W:=\begin{cases}\{({w},{v})\}&\text{if }\frac{1}{\gamma}=0,\\ \{(\tilde{w},\tilde{v})\ \lvert\ \tilde{w}\tilde{v}^{-1}=({w}{v}^{-1})^{t/t_{0}}\text{ and }[\tilde{w},\tilde{v}]_{(s_{0},r_{0})}<\infty\}&\text{else.}\end{cases}$ (5)
For the case $\frac{1}{\gamma}=0$ we do not introduce $({w}_{t},{v}_{t})$, but use the following convention: $\mathcal{M}$} is considered to be bounded from $L^{t}_{{w}_{t}}(\Omega)$ to $L^{t}_{{v}_{t}}(\Omega)$.
At first glance, the convention for $\frac{1}{\gamma}=0$ may appear unusual. However, it is specifically designed to ensure that the statements remain valid even in trivial cases, thereby eliminating the need for cumbersome case distinctions that could otherwise distract from the core arguments.
@@
If $\mathcal{M}$} is bounded from $L^{t}_{{w}_{t}}(\Omega)$ to $L^{t}_{{v}_{t}}(\Omega)$, then for all $f\in L^{p}_{w}(\Omega)$, $h\in L^{u}_{{v}^{-1}}(\Omega)$ there are weights $({w}_{0},{v}_{0})\in W$ such that $f\in L^{p_{0}}_{{w}_{0}}(\Omega),\ h\in L^{u_{0}}_{{v}_{0}^{-1}}(\Omega)$.
@@
$[{w}_{0},{v}_{0}]_{(s_{0},r_{0})}$ $\leq(\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t}_{{w}_{t}}\to L^{t}_{{v}_{t}}})^{\frac{t}{|\gamma|}}[{w},{v}]_{(s,r)}^{t/t_{0}},$ (6)
@@
$\|f\|_{L^{p_{0}}_{{w}_{0}}}\|h\|_{L^{u_{0}}_{{v}_{0}^{-1}}}\leq\kappa^{\frac{t}{|\gamma|}}\|f\|_{L^{p}_{w}}\|h\|_{L^{u}_{{v}^{-1}}}.$
@@
For $\frac{1}{\gamma}=0$, the assertion is trivial with $({w}_{0},{v}_{0}):=({w},{v})$. Thus, it suffices to assume $\gamma\in(0,\infty)$ by symmetry in $\gamma\leftrightarrow-\gamma$, $u_{0}\leftrightarrow p_{0}$, $r_{0}\leftrightarrow s_{0}$, $u\leftrightarrow p$, $s\leftrightarrow r$, ${w}_{0}\leftrightarrow{v}_{0}^{-1}$, and ${w}\leftrightarrow{v}^{-1}$. In particular, $p$ and $s$ are finite, $t_{0}=\frac{s_{0}}{\alpha}\in[1,\infty]$, and $t=\frac{s}{\alpha}\in[1,\infty)$. Let $f\in L^{p}_{w}(\Omega)$, $h\in L^{u}_{{v}^{-1}}(\Omega)$, and $\kappa\in(1,\infty)$. We assume that $f\neq 0$ (otherwise replace $f$ by some nonzero element in $L^{p}_{w}(\Omega)$). Define $H$ via $H^{t}{w}^{\alpha t}=|f|^{p}{w}^{p}$ and
$R:=\sum_{k=0}^{\infty}\frac{{\mathcal{M}}^{k}H}{(\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t}_{{w}^{\alpha}}\to L^{t}_{{v}^{\alpha}}})^{k}}.$
It follows immediately from this definition that
@@
    ${\mathcal{M}}R\leq\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t}_{{w}^{\alpha}}\to L^{t}_{{v}^{\alpha}}}R$, i.e. $R]_{(1,\infty)}\leq\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t}_{{w}^{\alpha}}\to L^{t}_{{v}^{\alpha}}}$,
@@
    $\|R\|_{L^{t}_{{v}^{\alpha}}}\leq\kappa\|H\|_{L^{t}_{{w}^{\alpha}}}$.
Let ${w}_{0}:=R^{-\frac{t}{\gamma}}{w}^{\frac{t}{t_{0}}}$ and ${v}_{0}:=R^{-\frac{t}{\gamma}}{v}^{\frac{t}{t_{0}}}$. Since $\frac{t\alpha}{\gamma}+\frac{t}{t_{0}t^{\prime}}=\frac{1}{t_{0}^{\prime}}$, Proposition 2.1 shows
$[{w}_{0}^{\alpha},{v}_{0}^{\alpha}]_{(t_{0},t_{0}^{\prime})}\leq[R^{-\frac{t\alpha}{\gamma}}]_{(\infty,\frac{\gamma}{t\alpha})}[{w}^{\frac{t\alpha}{t_{0}}},{v}^{\frac{t\alpha}{t_{0}}}]_{(t_{0},t_{0}t^{\prime}/t)}=[R]_{(1,\infty)}^{\frac{t\alpha}{\gamma}}[{w}^{\alpha},{v}^{\alpha}]_{(t,t^{\prime})}^{t/t_{0}}.$ (7)
Since ${w}_{0}^{\alpha},{v}_{0}^{\alpha}]_{(t_{0},t_{0}^{\prime})}=[{w}_{0},{v}_{0}]_{(s_{0},r_{0})}^{\alpha}$ and ${w}^{\alpha},{v}^{\alpha}]_{(t,t^{\prime})}=[{w},{v}]_{(s,r)}^{\alpha}$, estimate (6) follows from (7) via (ii). On $\{x\in\Omega\ \lvert\ |f(x)|>0\}$, (i) shows $|f|{w}_{0}\leq|f|H^{-\frac{t}{\gamma}}w^{\frac{t}{t_{0}}}=|f|^{\frac{p}{p_{0}}}{w}^{\frac{p}{p_{0}}}$, so that
$\|f{w}_{0}\|_{p_{0}}\leq\|f{w}\|_{p}^{\frac{p}{p_{0}}}.$
@@
$\|h{v}_{0}^{-1}\|_{u_{0}}\leq\|h{v}^{-1}\|_{u}\|{v}_{0}^{-1}{v}\|_{\gamma}=\|h{v}^{-1}\|_{u}\|R{v}^{\alpha}\|_{t}^{\frac{t}{\gamma}},$
so that the assertion follows since $\|R{v}^{\alpha}\|_{t}^{\frac{t}{\gamma}}\leq\kappa^{\frac{t}{\gamma}}\|H{w}^{\alpha}\|_{t}^{\frac{t}{\gamma}}=\kappa^{\frac{t}{\gamma}}\|f{w}\|_{p}^{1-\frac{p}{p_{0}}}$ by (iii). ∎
@@
    While Lemma 2.3 holds for $s,r\in(0,\infty]$, in applications it is often limited to $s,r\in(0,\infty)$. Clearly, if both $s=r=\infty$, then $1/\gamma=0$ and thus the assertion is trivial. If only one exponent is infinite, say $r=\infty$, then $t=1$, and hence $\mathcal{M}$} needs to be bounded from $L^{1}_{{w}^{s}}(\Omega)$ to $L^{1}_{{v}^{s}}(\Omega)$ with ${w}^{s},{v}^{s}]_{1}=[{w},{v}]_{(s,\infty)}^{s}<\infty$, which is typically not the case.
@@
If $\mathcal{M}$} is bounded from $L^{t}_{{w}_{t}}(\Omega)$ to $L^{t}_{{v}_{t}}(\Omega)$, then it holds
$L^{p}_{w}(\Omega)\subseteq\bigcup_{({w}_{0},{v}_{0})\in W}L^{p_{0}}_{{w}_{0}}(\Omega).$
@@
    $\mathcal{M}$} is bounded from $L^{t_{1j}}_{{w}_{t_{1j}}}(\Omega)$ to $L^{t_{1j}}_{{v}_{t_{1j}}}(\Omega)$ for all $j\in J$.
@@
    $T:\prod_{j=1}^{m}\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}S^{-1}(L^{p_{0j}}_{{w}_{0j}}(\Omega))\to L^{0}(\Omega),$
    and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all vectors of weights $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}:=W_{1}\times\cdots\times W_{m}$ and for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{p_{0j}}_{{w}_{0j}}(\Omega))$ it holds
    $\|T\vec{f}\|_{L^{q_{0}}_{{v}_{0}}}\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{w}_{0j}}}.$ (9)
Then $T\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{p_{1j}}_{{w}_{1j}}(\Omega))$, and for all $\kappa\in(1,\infty)$ it holds
$\|T\vec{f}\|_{L^{q_{1}}_{{v}_{1}}}\leq\kappa^{\beta}\phi(\vec{C}_{\kappa})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{1j}}_{{w}_{1j}}},$ (10)
where $\beta:=\sum_{j\in J}\frac{t_{1j}}{|\gamma_{j}|}$ and $C_{\kappa j}:=(\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t_{1j}}_{{w}_{t_{1j}}}\to L^{t_{1j}}_{{v}_{t_{1j}}}})^{\frac{t_{1j}}{|\gamma_{j}|}}[{w}_{1j},{v}_{1j}]_{(s_{1j},r_{1j})}^{t_{1j}/t_{0j}}$.

### v2

@@ 1 Introduction
As a matter of fact, this symmetry also shows its benefits for the extrapolation results. The theory of $L^{p}$-extrapolation has a long history starting with the works by García-Cuerva and Rubio de Francia [8, 17], who realized that boundedness of an operator $T$ on all Muckenhoupt weighted $L^{p}$-spaces implies boundedness on all Muckenhoupt weighted $L^{q}$-spaces for the range $q\in(1,\infty)$. They also showed that vector-valued estimates can be extracted from scalar-valued weighted estimates. Later, Buckley [2] found quantitative bounds for the maximal operator in Muckenhoupt weighted spaces in terms of the weight characteristic. Restricting the weight classes for the initial spaces of extrapolation, Auscher and Martell proved first versions of limited-range extrapolation in [1]. For a detailed history of the theory, we refer to Nieraeth [13], where a comprehensive theory for multilinear extrapolation is established. In particular, in the multilinear case $T:\prod_{j=1}^{m}L^{p_{j}}_{{w}_{j}}\to L^{p}_{{w}}$ with $\frac{1}{p}=\sum_{j=1}^{m}\frac{1}{p_{j}}$ some of the $p_{j}$’s (but not all of them) are allowed to be infinity in the target spaces of the extrapolation. In [11], the authors provided conditions that ensure extrapolation towards $p=\infty$. In the linear case, they also showed that this is possible in the off-diagonal setting, that is when $p\neq p_{1}$. Explicit bounds for the results of [11] were found [3]. Recently, Nieraeth [14] found extrapolation results from Lebesgue spaces over $\sigma$-finite measure to general quasi-Banach function spaces, which includes mixed-norm estimates. Our contribution to the theory of extrapolation in weighted spaces is as follows:
@@
    We extend the results of [3] and [11]: We include the range $(0,1]$ for all initial and target spaces of the extrapolation. We show off-diagonal estimates even in the multilinear case and allow for mixed $L^{\mathbf{p}}$-norms and two-weight estimates. We provide explicit bounds as in [3], which are sharp in all linear and scalar-valued cases considered in [11]. In comparison to [3] and [11] our proofs are significantly shorter and leaner. This is achieved by systematically exploiting the symmetry of the weight classes. In fact, Theorem 2.6 contains Theorem 2.3 in [11], where the authors are forced to consider three different cases, by specialising to the case $p_{0}\in[1,\infty]$, $p\in(1,\infty]$ and $r_{0}:=p_{0}^{\prime}$, $r=p^{\prime}$. We warn the reader that our $s$ is called $r$ in [11]. Finally, we work in the setting of general $\sigma$-finite measure spaces as in [14] instead of $\mathbb{R}^{d}$.
@@
We fix $(\Omega,\mu)$ together with a basis of sets $U$ for the rest of this article and write $L^{0}(\Omega;X)$ for the space of strongly measurable functions, where $X$ is a quasi-Banach space. In the scalar-valued case, we simply write $L^{0}(\Omega)$.
@@
Note that ${\mathcal{M}}f$ is the countable supremum of measurable functions and hence measurable. Moreover, for $f\in L^{1}(\Omega)$ we have ${\mathcal{M}}f>0$ whenever $f\neq 0$, see [14, Proposition 2.16].
@@
    $\|{\mathcal{M}}\|_{L^{t}_{{v}}(\Omega)\to L^{t}_{{v}}(\Omega)}\leq c(t)[{v}]_{t}^{t^{\prime}}\quad\text{for all $t\in(1,\infty]$ and weights ${v}$}.$
@@
    As observed in [14, Remark 4.8], the formulation in Theorem 1.2 implies a result in terms of extrapolation pairs. Indeed, if
@@
    then the choice $V:=\mathcal{F}$, $S(f,g):=f$, and $T(f,g):=g$ shows that a bound $\|g\|_{L^{q_{0}}_{{w}_{0}}(\Omega)}\leq\phi([{w}]_{(s_{0},r_{0})})\|f\|_{L^{p_{0}}_{{w}_{0}}(\Omega)}$ for all $(f,g)\in\mathcal{F}$ and weights ${w}_{0}$ with ${w}_{0}]_{(s_{0},r_{0})}<\infty$ extrapolates to a bound $\|g\|_{L^{q}_{{w}}(\Omega)}\leq\phi_{s_{0},r_{0},s}([{w}]_{(s,r)})\|f\|_{L^{p}_{{w}}(\Omega)}$ for all $(f,g)\in\mathcal{F}$ and weights w{w} with ${w}]_{(s,r)}<\infty$.
@@
    We can deduce a bound $L^{1}_{w}(\mathbb{R}^{d})\to L^{\infty}_{w}(\mathbb{R}^{d})$ for all weights w{w} with ${w}]_{2}<\infty$ whenever we already know a bound $L^{\frac{2}{3}}_{{w}_{0}}(\mathbb{R}^{d})\to L^{2}_{{w}_{0}}(\mathbb{R}^{d})$ for all weights ${w}_{0}$ with ${w}_{0}]_{1}<\infty$. Indeed, setting
@@
    we obtain this assertion as a particular instance of Theorem 1.1 since $\|{\mathcal{M}}\|_{L^{t}_{{w}}(\mathbb{R}^{d})\to L^{t}_{{w}}(\mathbb{R}^{d})}\leq c(t)[{w}]_{t}^{t^{\prime}}$ for all $t\in(1,\infty]$ and weights w{w}. This shows that by detaching the exponents for the weight classes from the exponents for the initial and target spaces for the extrapolation, Theorem 1.1 can cover situations that can not be treated by [11, Theorem 2.3], where $p$ has to be strictly larger than $1$.
@@ 2 Limited-Range Multilinear Off-Diagonal Extrapolation
$W:=\begin{cases}\{({w},{v})\}&\text{if }\frac{1}{\gamma}=0,\\ \{(\tilde{w},\tilde{v})\ \lvert\ \tilde{w}\tilde{v}^{-1}={w}{v}^{-1}\text{ and }[\tilde{w},\tilde{v}]_{(s_{0},r_{0})}<\infty\}&\text{else.}\end{cases}$ (5)
For the case $\frac{1}{\gamma}=0$ we do not introduce $({w}_{t},{v}_{t})$, but use the following convention: $\mathcal{M}$} is considered to be bounded from $L^{t}_{{v}_{t}}(\Omega)$ to $L^{t}_{{w}_{t}}(\Omega)$.
At first glance, the convention for $\frac{1}{\gamma}=0$ may appear unusual. However, it is specifically designed to ensure that the statements remain valid even in trivial cases, thereby eliminating the need for cumbersome case distinctions that could otherwise distract from the core arguments. Moreover, if w=v{w}={v}, the definition of the weight class $W$ ensures that ${w}_{0}={v}_{0}$ for $({w}_{0},{v}_{0})\in W$.
@@
If $\mathcal{M}$} is bounded from $L^{t}_{{v}_{t}}(\Omega)$ to $L^{t}_{{w}_{t}}(\Omega)$, then for all $f\in L^{p}_{v}(\Omega)$, $h\in L^{u}_{{w}^{-1}}(\Omega)$ there are weights $({w}_{0},{v}_{0})\in W$ such that $f\in L^{p_{0}}_{{v}_{0}}(\Omega),\ h\in L^{u_{0}}_{{w}_{0}^{-1}}(\Omega)$.
@@
$[{w}_{0},{v}_{0}]_{(s_{0},r_{0})}$ $\leq(\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t}_{{v}_{t}}\to L^{t}_{{w}_{t}}})^{\frac{t}{|\gamma|}}[{w},{v}]_{(s,r)}^{t/t_{0}},$ (6)
@@
$\|f\|_{L^{p_{0}}_{{v}_{0}}}\|h\|_{L^{u_{0}}_{{w}_{0}^{-1}}}\leq\kappa^{\frac{t}{|\gamma|}}\|f\|_{L^{p}_{v}}\|h\|_{L^{u}_{{w}^{-1}}}.$
@@
For $\frac{1}{\gamma}=0$, the assertion is trivial with $({w}_{0},{v}_{0}):=({w},{v})$. Thus, it suffices to assume $\gamma\in(0,\infty)$ by symmetry in $\gamma\leftrightarrow-\gamma$, $u_{0}\leftrightarrow p_{0}$, $r_{0}\leftrightarrow s_{0}$, $u\leftrightarrow p$, $s\leftrightarrow r$, ${w}_{0}\leftrightarrow{v}_{0}^{-1}$, and ${w}\leftrightarrow{v}^{-1}$. Moreover, replacing (w,v)({w},{v}) by $({w}^{\alpha},{v}^{\alpha})$ if necessary, we may assume $\alpha=1$ by Proposition 2.1(ii). In particular, $p$ is finite, $s=t\in[1,\infty)$ and $s_{0}=t_{0}\in[1,\infty]$ with $s^{\prime}=r$ and $s_{0}^{\prime}=r_{0}$. Let $f\in L^{p}_{w}(\Omega)$, $h\in L^{u}_{{v}^{-1}}(\Omega)$, and $\kappa\in(1,\infty)$. We assume that $f\neq 0$ (otherwise replace $f$ by some nonzero element in $L^{p}_{w}(\Omega)$). Define $H$ via $H^{s}{w}^{s}=|f|^{p}{v}^{p}$ and with the notation ${\mathcal{M}}_{t}g:={\mathcal{M}}g_{t}$, where $g_{t}:=g{w}{v}^{-1}$,
$R:=\sum_{k=0}^{\infty}\frac{{\mathcal{M}}_{t}^{k}H}{(\kappa^{\prime}\|{\mathcal{M}}_{t}\|_{L^{s}_{{w}}\to L^{s}_{{w}}})^{k}}.$
Since $\|{\mathcal{M}}_{t}\|_{L^{s}_{{w}}\to L^{s}_{{w}}}=\|{\mathcal{M}}\|_{L^{s}_{{v}}\to L^{s}_{{w}}}$, it follows from this definition and Proposition 2.1(iv) that
@@
    ${\mathcal{M}}R_{t}\leq\kappa^{\prime}\|{\mathcal{M}}\|_{L^{s}_{{v}}\to L^{s}_{{w}}}R$ for $R_{t}:=R{w}{v}^{-1}$, i.e. $R_{t},R]_{(1,\infty)}\leq\kappa^{\prime}\|{\mathcal{M}}\|_{L^{s}_{{v}}\to L^{s}_{{w}}}$,
@@
    $\|R\|_{L^{s}_{{w}}}\leq\kappa\|H\|_{L^{s}_{{w}}}$.
Let ${w}_{0}:=R^{-\frac{s}{\gamma}}{w}^{\frac{s}{s_{0}}}$ and ${v}_{0}:=R_{t}^{-\frac{s}{\gamma}}{v}^{\frac{s}{s_{0}}}$, so that ${w}_{0}{v}_{0}^{-1}={w}{v}^{-1}$ in virtue of $\frac{s}{\gamma}+\frac{s}{s_{0}}=1$. Since $\frac{s}{\gamma}+\frac{s}{s_{0}r}=\frac{1}{r_{0}}$, Proposition 2.1 shows
$[{w}_{0},{v}_{0}]_{(s_{0},r_{0})}\leq[R^{-\frac{s}{\gamma}},R_{t}^{-\frac{s}{\gamma}}]_{(\infty,\frac{\gamma}{s})}[{w}^{\frac{s}{s_{0}}},{v}^{\frac{s}{s_{0}}}]_{(s_{0},s_{0}r/s)}=[R_{t},R]_{(1,\infty)}^{\frac{s}{\gamma}}[{w},{v}]_{(s,r)}^{\frac{s}{s_{0}}}.$ (7)
Thus, (6) follows from (7) via (ii). On $\{x\in\Omega\ \lvert\ |f(x)|>0\}$, (i) shows $|f|{v}_{0}\leq|f|(H{w}{v}^{-1})^{-\frac{s}{\gamma}}{v}^{\frac{s}{s_{0}}}=(|f|{v})^{\frac{p}{p_{0}}}$, so that
$\|f{v}_{0}\|_{p_{0}}\leq\|f{v}\|_{p}^{\frac{p}{p_{0}}}.$
@@
$\|h{w}_{0}^{-1}\|_{u_{0}}\leq\|h{w}^{-1}\|_{u}\|{w}_{0}^{-1}{w}\|_{\gamma}=\|h{w}^{-1}\|_{u}\|R{w}\|_{s}^{\frac{s}{\gamma}},$
so that the assertion follows since $\|R{w}\|_{s}^{\frac{s}{\gamma}}\leq\kappa^{\frac{s}{\gamma}}\|H{w}\|_{s}^{\frac{s}{\gamma}}=\kappa^{\frac{s}{\gamma}}\|f{v}\|_{p}^{1-\frac{p}{p_{0}}}$ by (iii). ∎
@@
    While Lemma 2.3 holds for $s,r\in(0,\infty]$, in applications it is often limited to $s,r\in(0,\infty)$. Clearly, if both $s=r=\infty$, then $1/\gamma=0$ and thus the assertion is trivial. If only one exponent is infinite, say $r=\infty$, then $t=1$, and hence $\mathcal{M}$} needs to be bounded from $L^{1}_{{v}^{s}}(\Omega)$ to $L^{1}_{{w}^{s}}(\Omega)$ with ${w}^{s},{v}^{s}]_{1}=[{w},{v}]_{(s,\infty)}^{s}<\infty$, which is typically not the case.
@@
If $\mathcal{M}$} is bounded from $L^{t}_{{v}_{t}}(\Omega)$ to $L^{t}_{{w}_{t}}(\Omega)$, then it holds
$L^{p}_{v}(\Omega)\subseteq\bigcup_{({w}_{0},{v}_{0})\in W}L^{p_{0}}_{{v}_{0}}(\Omega).$
@@
    $\mathcal{M}$} is bounded from $L^{t_{1j}}_{{v}_{t_{1j}}}(\Omega)$ to $L^{t_{1j}}_{{w}_{t_{1j}}}(\Omega)$ for all $j\in J$.
@@
    $T:\prod_{j=1}^{m}\bigcup_{({w}_{0j},{v}_{0j})\in W_{j}}S^{-1}(L^{p_{0j}}_{{v}_{0j}}(\Omega))\to L^{0}(\Omega),$
    and a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ increasing in each component such that for all vectors of weights $(\vec{{w}}_{0},\vec{{v}}_{0})\in\vec{W}:=W_{1}\times\cdots\times W_{m}$ and for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{p_{0j}}_{{v}_{0j}}(\Omega))$ it holds
    $\|T\vec{f}\|_{L^{q_{0}}_{{w}_{0}}}\leq\phi([\vec{{w}}_{0},\vec{{v}}_{0}]_{(\vec{s}_{0},\vec{r}_{0})})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{0j}}_{{v}_{0j}}}.$ (9)
Then $T\vec{f}$ is well-defined for all $\vec{f}\in\prod_{j\in J}S_{j}^{-1}(L^{p_{1j}}_{{v}_{1j}}(\Omega))$, and for all $\kappa\in(1,\infty)$ it holds
$\|T\vec{f}\|_{L^{q_{1}}_{{w}_{1}}}\leq\kappa^{\beta}\phi(\vec{C}_{\kappa})\prod_{j\in J}\|S_{j}f_{j}\|_{L^{p_{1j}}_{{v}_{1j}}},$ (10)
where $\beta:=\sum_{j\in J}\frac{t_{1j}}{|\gamma_{j}|}$ and $C_{\kappa j}:=(\kappa^{\prime}\|{\mathcal{M}}\|_{L^{t_{1j}}_{{v}_{t_{1j}}}\to L^{t_{1j}}_{{w}_{t_{1j}}}})^{\frac{t_{1j}}{|\gamma_{j}|}}[{w}_{1j},{v}_{1j}]_{(s_{1j},r_{1j})}^{t_{1j}/t_{0j}}$.